OpenStudy (anonymous):

ow can u rewrite these logarithmic equations an equivalent exponent equation or the other way around m=loga x is equivalent to am=x1. log3 5=y,2. loga 7=-2then solvelog2x=-3,logx9=1/2

OpenStudy (anonymous):

need help solving heres attachment

OpenStudy (anonymous):

someone was helping then we got disconnected

OpenStudy (anonymous):

You were given this lograithmic form. \[m=log_{a}x\] And your question tells you rewrite in a different form. \[y=log_{3}5\] Tell me, using the logarithmic form that was givne to you, what is m=?, a=? and x=? in your first question.

OpenStudy (anonymous):

no how to rewrite them now dont know how to solve them need help

OpenStudy (anonymous):

3y=5 and 3base a-2=7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

It doesn't say solve.

OpenStudy (anonymous):

That's for part b...

OpenStudy (anonymous):

i did the first part not sure if correct

OpenStudy (anonymous):

Don't just look at a question and instantly think you have to solve it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Actually read the question. It only says ot rewrite it. And That's what I was asking you before.

OpenStudy (anonymous):

Read what I wrote before and tell me what m=?, a=? and x=?

OpenStudy (anonymous):

m=log a=5y x=3y=5

OpenStudy (anonymous):

m=log. What?!?!?...

OpenStudy (anonymous):

ok, i guess i am wrong

OpenStudy (anonymous):

The very start of your worksheet says "The logarithmic form \[m=log_{a}x\] can be rewritten as \[a^m=x\] I don't see any difficulty in matching up your numbers and values with that form and converting it into that form...

OpenStudy (anonymous):

You're up to this stage/level in mathematics, but you're still not sure how to match up numbers and values with the positions they were in and then using that to convert it into that speicif form, which was told to you at the very beginning.

OpenStudy (anonymous):

i never did anything like this before and been out of school in a while.

OpenStudy (anonymous):

This is just playing match ups.

OpenStudy (anonymous):

You match this to that. And that to this. What's so hard about that? All I see is someone who is just not trying. I will match up the first one and then you can do the rest. This is so simpple. In that equation \[m=log_{a}x\] compared to your very first question \[y=log_{3}5\] I see that m=y. Now tell me what a=? and what x=?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Dude, what does that mean? You haven't stated anything nor have you answered my question. You just replied with the number 3 when I told you to give me 2 answers. Don't just reply back with "3". That's someone who's simply being lazy.

OpenStudy (anonymous):

\[y=log_{3}5\] \[3^y=5\]